A hydrogen atom in the ground state absorbs a 20 eV photon. What is the speed of the liberated electron?
Answer #1

ground state energy = -13.6 eV

Thus 13.6 eV is used to get out of the atom.

Remaining energy = 6.4 eV

v = sqrt(2K/m) = 1.5X10^6 m/s
Answer #2

apply KE = PE

0.5 mv^2 = eV

V^2 = 2*1.6*10^-19 * 20/9.11*10-^-31

V = 2.65*10^6 m/s